OpenStudy (anonymous):

Determine the value of r so the line passing through each pair of points has the given slope.

OpenStudy (anonymous):

(-5,5) (-3,r) m=2/3

OpenStudy (anonymous):

><''

OpenStudy (anonymous):

i would totally help you, but i have no clue how to do that.

OpenStudy (anonymous):

I know. :P

OpenStudy (anonymous):

D:

OpenStudy (anonymous):

the slope is \[\frac{r-5}{-3-(-5)}=\frac{r-5}{2}\] set \[\frac{r-5}{2}=\frac{2}{3}\] and solve for \(r\)

OpenStudy (anonymous):

clear how to do that?

OpenStudy (anonymous):

...

OpenStudy (anonymous):

Sort of..

OpenStudy (anonymous):

But I got 17/3...which i'm sure is incorrect.

OpenStudy (anonymous):

lets see

OpenStudy (anonymous):

start with \[3(r-5)=2\times 2\] \[3r-15=4\] \[3r=19\] \[r=\frac{19}{3}\] you were close

OpenStudy (anonymous):

...hmm....

OpenStudy (anonymous):

ohh.. I see.

OpenStudy (anonymous):

Thanks. c;